Video Planet (VP) sells a big screen TV package consisting of a 60-inch plasma TV, a universal remote, and on-site installation by VP staff. The installation includes programming the remote to have the TV interface with other parts of the customer’s home entertainment system. VP concludes that the TV, remote, and installation service are separate performance obligations. VP sells the 60-inch TV separately for $1,910 and sells the remote separately for $180, and offers the entire package for $2,220. VP does not sell the installation service separately. VP is aware that other similar vendors charge $230 for the installation service. VP also estimates that it incurs approximately $180 of compensation and other costs for VP staff to provide the installation service. VP typically charges 50% above cost on similar sales.

Required:
1. to 3. Calculate the stand-alone selling price of the installation service using each of the following approaches.
Answer #1

Price will be as follows:

Adjusted market assessment = $230 i.e. the price charged by similar vendors

Expected Cost plus margin = 180 + 180*50% = $270

Residual = Total package revenue – Other products

= 2,220-1,910-180

= $130